# Identifying Spurious Correlations using Counterfactual Alignment

### v1

@@ 4.4 Validation by inducing spurious correlations
(a) CF for smiling showing eyebrows are unchanged. The horizontal line indicates the prediction of arched_eyebrows is not influenced by the features used for smiling in this image.
@@
(b) CF for the modified smiling classifier with an arched eyebrow spurious correlation showing that eyebrows are now changed.
@@
Figure 5: Example of detecting a spurious correlation in a biased classifier. The classifier is biased with arched eyebrows and this is observed in the alignment plot as well as in the counterfactual image. In order to decrease the prediction of the (now biased) classifier, the CF image can remove the arched eyebrows as well.
4.5 Rectifying bias by composing classifiers
@@
Correcting model bias and spurious correlations is an active area of research. Approaches include averaging model weights [36] or methods of invariant optimization [26, 16, 24, 37]. Composing classifiers can change their response to specific samples that contain relevant features as shown in §4.4. By using this additive composition approach we avoid the variance of model training which makes it easy and reproducible to perform experiments.
@@
Next, this unbiasing is scaled up to a larger number of samples and classifiers. A collection of classifiers that contain varying spurious relationships to a big_nose base classifier (shown in Fig. 7a) are modified to remove their bias until their relative change is as close to 0 as possible. This experiment is performed with a train, validation, and hold out test set in order to demonstrate the generalization of this unbiasing process to unseen data.
@@
The optimization to compute $\beta$ is performed using a pseudo gradient descent where the gradient is approximated by the mean relative change between the target classifier and the big_nose classifier. By subtracting the relative change from $\beta$, scaled by a learning rate, the relative change ($\psi$) with that classifier will be reduced. All together our training objective (including momentum) is
@@
The resulting training metrics in Fig. 7 show model biases being summarized and minimized similar to a differentiable training loss. We observe a bias generalization gap between the train, valid, and test sets, indicating some degree of overfitting. Early stopping (red vertical line) is used on the validation set to determine the optimal parameters. We use momentum during training to average over the noise. The resulting reduction in bias is shown in Fig 7b. All biases are considerably reduced except for attractive and big_nose, which could be the result of overlapping features used by both classifiers that cannot be changed.
5 Limitations
There are limitations and challenges to CF generation [35] where bias can also exist in the CF generation method. Furthermore, although we are optimistic that our method will generalize to additional domains, we focus our current analysis on face attribute classification. Although we do not demonstrate the generalizability of our method to additional domains, using face attribute classification is a deliberate choice enabling qualitative evaluation, in addition to our quantitative evaluation. Additionally, since CF generation relies on a separate autoencoder, improving the representational capability of the autoencoder may improve the fidelity of the CF generation.
6 Conclusion
In this work we propose the counterfactual (CF) alignment approach to reason about the feature relationships in classifiers in aggregate statistics, to locate specific examples with spurious correlations, and propose strategies to mitigate the effect of spurious correlations. These claims are supported with an analysis of face attribute classifiers that identify many expected and unexpected spurious correlations such as attractive with heavy_makeup and heavy_makeup with big_nose. The validity of the CF alignment method is confirmed by inducing and quantifying spurious correlations via additive composition. We then explore rectifying classifiers based on observed spurious correlations. Overall, the proposed approach may serve as an end-to-end or human-in-the-loop system to automatically detect, quantify, and correct spurious correlations for image classification tasks that lead to biased classifier outputs.
7 Acknowledgments
Supplementary Material
8 Appendix
@@
Figure 10: Visualization of the training from §4.5 reducing the bias of big_nose. The model being evaluated is $f^{\prime}_{target}(x)=f_{target}(x)+\beta f_{big\_nose}(x)$ where $\beta$ is specified by the coefficient column. The metric here is relative change computed between the $f^{\prime}_{target}(x)$, where $target$ is specified in the caption, and the classifier specified by the column header. The top row is from the training set performed on batches (which explains the variance) and the second row is the results on the entire validation set. The row of each table is for a specific iteration. The iterations proceed downward. The training of attractive shown here does not result in a debiased model, the relative change diverges from 0 indicating a failed training.

### v2

@@ 4.4 Validation by inducing spurious correlations
(a) CF for the classifier smiling showing eyebrows are unchanged. The horizontal line indicates the prediction of arched_eyebrows is not influenced by the features used for smiling in this image.
@@
(b) CF for a modified smiling classifier which has been combined with an arched_eyebrows classifier.
@@
Figure 5: Example of detecting a spurious correlation in a biased classifier. The classifier is biased with arched eyebrows and this is observed in the alignment plot as well as in the counterfactual image. The relative change is now 0.97 compared to 0.01 for the unchanged smiling classifier.
4.5 Analysis of robust optimization methods
An application of CF alignment is to assess the effectiveness of bias mitigation techniques. Various biased training settings have been developed to induce spurious correlations with sensitive attributes in models and then employ bias mitigation strategies to counter them. The experiments we present are designed to demonstrate that measuring the CF alignment with sensitive attributes can measure the impact of these methods.
To properly compare these methods, experiments are grouped based on the specific training configuration used so each baseline is unique to that setting. For this reason, these models cannot be compared side by side for the same dataset.
The bias mitigation methods we use are selected because their authors make their code available to generate pre-trained models that can be studied post-hoc. Using only pre-trained models further demonstrates the utility of our proposed approach where our method is versatile enough to evaluate these existing pre-trained models.
Experiments in this section are performed on the held out test datasets and selected to be balanced such that there is a balanced distribution of positive and negative examples with the sensitive attribute (e.g. in the CelebA dataset, 1/4 of the samples are labeled to have Blond Hair and be Male, 1/4 Blond Hair and be Not Male, 1/4 Not Blond Hair and be Not Male, and 1/4 Not Blond Hair and be Male).
During classification experiments, the samples evaluated are selected to be inversely correlated with the sensitive attribute to cause the most discrepancy in performance (e.g. Samples are selected that are labeled Blond Hair and Male as well as Not Blond Hair and Not Male).
4.5.1 GroupDRO on Waterbirds
Work on GroupDRO Sagawa et al. (2020) constructed the Waterbirds task such that there is a spurious correlation between birds (from the CUB dataset (Wah et al., 2011)) with backgrounds that contain water or land.
Three models are evaluated for this experiment, a baseline model which was induced to rely on spurious correlations, a model trained with Group Distributionally Robust Optimization (DRO) (Sagawa et al., 2020) which aims to minimize error over groups (which are known during training, in this case the background class), a model using Just Train Twice (JTT) (Liu et al., 2021) which uses a two stage approach that boosts misclassified training examples in a second training cycle. The classifiers that these models are aligned with are trained to predict if the background of an image is land or water using the same training dataset labels from the Waterbirds dataset.
A limitation of this analysis is the limited latent variable model used (A VQ-GAN trained on Open Images (Krasin et al., 2017)) that has difficulty modulating features (likely because its training domain is broad). A model that is better at representing birds would be able to generate better counterfactuals that are easier to interpret.
In Figure 6 two models are evaluated on the same image illustrating a difference in CF alignment results to a classifier that predicts the background. The relative change with the background classifier is much higher for the baseline model than the model trained with DRO (0.35 vs 0.25) indicating that training with GroupDRO prevented the models reliance on features associated with the background classifier. A magnified view of the counterfactuals of the bird’s head indicate nostril size being reduced and colors becoming brighter in the model trained with DRO. This may indicate larger nostril sizes are associated with water birds which aligns with the family of waterbirds, Procellariiformes, having an enlarged nasal gland at the base of the beak for secreting salt water (Ehrlich et al., 1988). While the mallard in the image is not a Procellariiforme, nostril size may be a feature used by the classifier in general as many in the waterbirds are seabirds.
Aggregate results over 4097 samples from the test set of the waterbirds dataset reveals a reduction in mean relative change using DRO shown in Table 1. This indicates that the features used by the baseline classifier overlap the land and water classifier and these spurious correlations were reduced when utilizing DRO. We observe that this improved model, with a lower relative change, achieves a higher classification accuracy.
(a) Evaluating a baseline model trained that is expected to use spurious correlations.
(b) Evaluating a model trained using Distributionally Robust Optimization (DRO) to reduce the spurious correlation with the background.
(c) A magnified view of the birds head which shows differences in nostril size and head color.
Figure 6: Counterfactuals generated for models trained on the waterbirds dataset together with CF alignment plots. (a) and (b) The relative change between the waterbird classifier and the background classifier is shown in the lower right of the plot. (c) The nostril size is reduced in the counterfactual for the DRO model indicating that a larger nostril size is associated with waterbirds.
Target
Relative change $\downarrow$
w/water classifier
Relative change $\downarrow$
w/land classifier
Classification $\uparrow$
AUC
Waterbird (baseline) 0.35±0.02 -0.36±0.03 0.69
Waterbird (groupdro) $0$0.22±0.03 0.92
Waterbird (jtt) 0.42±0.03 -0.43±0.00 0.65
Table 1: Aggregate relative change metrics for models trained on the waterbirds test dataset (N=4097). A reduction in relative change with the land and water background classifiers indicates DRO has reduced the spurious correlation.
4.5.2 GroupDRO on CelebA
GroupDRO models are also evaluated on the CelebA dataset on the task of Blond_Hair and the sensitive attribute Male. JTT is not evaluated on this task because we were unable to generate pre-trained weights using the provided code due to memory issues that we were unable to resolve.
In Table 2 we can see the relationship between Blond_Hair and Male is negative in the baseline and then is reduced by the GroupDRO training and becomes slightly positive. This reduction results in a higher classification accuracy on the CelebA test set.
Target
Relative change $\downarrow$
w/male classifier
Classification $\uparrow$
AUC
Blond_Hair (baseline) -0.052±0.01 0.91
Blond_Hair (groupdro) 0.025±0.00 0.95
Table 2: Methods evaluated on models that predict Blond_Hair trained on the CelebA dataset. Evaluations are performed on 1024 samples from the test set.
4.5.3 FLAC on CelebA
The work Fairness-Aware Representation Learning by Suppressing Attribute-Class Associations (FLAC) (Sarridis et al., 2023) uses pairwise similarity between the between the model representation and the representation of a bias-capturing classifier that predicts the sensitive attribute. The bias capturing classifier acts as a proxy for labels and provides embeddings that can be used to regularize the feature space of the classifier we are training. Their approach is to minimize the difference, in feature space, between samples with different sensitive attributes and the same target label as well as increase the difference between samples with the same sensitive attributes and the different target label. In Table 3 the relative change is shown to be reduced and the classification performance improves.
Target Alpha
Relative change $\downarrow$
w/male classifier
Classification $\uparrow$
AUC
Blond_Hair 0 -0.081±0.016 0.77
Blond_Hair 30000 -0.039±0.009 0.93
Blond_Hair 60000 -0.051±0.010 0.92
Table 3: Aggregate relative change of models trained with FLAC on 1024 samples from the CelebA test set. The strength of this regularization is controlled with an $\alpha$ parameter.
5 Limitations
There are limitations and challenges to CF generation Verma et al. (2020) where bias can also exist in the CF generation method. Furthermore, although we are optimistic that our method will generalize to additional domains, we focus our current analysis on face-attribute classification. Although we do not demonstrate the generalizability of our method to additional domains, using face-attribute classification is a deliberate choice enabling qualitative evaluation, in addition to our quantitative evaluation. Additionally, since CF generation relies on a separate autoencoder, improving the representational capability of the autoencoder may improve the fidelity of the CF generation.
6 Conclusion
In this work we propose counterfactual (CF) alignment along with the relative change metric §3. We demonstrate that this method enables us to reason about the feature relationships between classifiers in aggregate and to locate specific examples where the spurious correlations are used. These claims are supported with an analysis of face-attribute classifiers that identify expected and unexpected spurious correlations.
We observe that if correlated attributes are presented to the classifier, this does not cause feature usage to be correlated §4.1. We also observe that if uncorrelated attributes are presented to the classifier, the classifier may still construct spurious feature relationships.
The validity of the CF alignment method is confirmed by inducing and quantifying spurious correlations via additive composition §4.4. Classifiers are composed together to create a classifier with a known spurious correlation and then this is observed using CF alignment.
We then explore classifiers trained using robust optimization methods to demonstrate the applicability to black box classifiers and to provide more visibility into what these methods achieve §4.5.
Overall, the proposed approach may serve as an end-to-end or human-in-the-loop system to automatically detect, quantify, and correct spurious correlations for image classification tasks that lead to biased classifier outputs.
7 Acknowledgments
Appendix A Appendix
A.1 Rectifying bias by composing classifiers
@@
This section just serves as an example to better understand a use of the CF alignment idea. We don’t claim this method is competitive to other methods which correct model bias and spurious correlations. Related approaches include averaging model weights Wortsman et al. (2022) or methods of invariant optimization Sagawa et al. (2020); Krueger et al. (2021); Rieger et al. (2020); Zeng et al. (2023) and approaches for balanced dataset sampling Singh et al. (2021). Composing classifiers can change their response to specific samples that contain relevant features as shown in §4.4. By using this additive composition approach we avoid the variance of model training which makes it easy and reproducible to perform experiments.
@@
Next, this unbiasing is scaled up to a larger number of samples and classifiers. A collection of 12 classifiers that contain varying spurious relationships to a big_nose base classifier (shown in Fig. 9a) are modified to remove their bias until their relative change is as close to 0 as possible. This experiment is performed with a train, validation, and hold out test set in order to demonstrate the generalization of this unbiasing process to unseen data.
@@
Optimization to compute each $\beta$ is performed using a pseudo gradient descent where the gradient is approximated by the mean relative change between the target classifier and the big_nose classifier. By subtracting the relative change from $\beta$, scaled by a learning rate, the relative change ($\psi$) with that classifier will be reduced. All together our training objective (including momentum) is
@@
The resulting training metrics in Fig. 9 show model biases being minimized similar to a differentiable training loss. Looking at the mean relative change allows us to summarize the bias of the model with respect to the big_nose classifier who’s relationship we aim to remove. We observe a bias generalization gap between the train, valid, and test sets, indicating some degree of overfitting. Early stopping (red vertical line) is used on the validation set to determine the optimal parameters. We use momentum during training to average over the noise.
The resulting reduction in bias is shown per classifier before and after training in Fig 10. 10 of the 12 classifiers being optimized have the bias considerably reduced (closer to 0) except for attractive and big_nose, which could be the result of overlapping features used by both classifiers that cannot be changed.
@@
Figure 11: Visualization of the training from §A.1 reducing the bias of big_nose. The model being evaluated is $f^{\prime}_{target}(x)=f_{target}(x)+\beta f_{big\_nose}(x)$ where $\beta$ is specified by the coefficient column. The metric here is relative change computed between the $f^{\prime}_{target}(x)$, where $target$ is specified in the caption, and the classifier specified by the column header. The top row is from the training set performed on batches (which explains the variance) and the second row is the results on the entire validation set. The row of each table is for a specific iteration. The iterations proceed downward. The training of attractive shown here does not result in a debiased model, the relative change diverges from 0 indicating a failed training.